A researcher studied the relation between psychotherapists degree of empathy and their patients satisfaction with therapy.Pair Number Patient Satisfaction Therapist Empathy 74 6 2 90 7 3 31 5 4 47 3
Answer #1
X Y XY
74 6 444 5476 36
90 7 630 8100 49
31 5 155 961 25
47 3 141 2209 9
Ʃx = 242
Ʃy = 21
Ʃxy = 1370
Ʃx² = 16746
Ʃy² = 119
Sample size, n = 4
x̅ = Ʃx/n = 242/4 = 60.5
y̅ = Ʃy/n = 21/4 = 5.25
SSxx = Ʃx² - (Ʃx)²/n = 16746 - (242)²/4 = 2105
SSyy = Ʃy² - (Ʃy)²/n = 119 - (21)²/4 = 8.75
SSxy = Ʃxy - (Ʃx)(Ʃy)/n = 1370 - (242)(21)/4 = 99.5

c)

Correlation coefficient, r = SSxy/√(SSxx*SSyy) = 99.5/√(2105*8.75) = 0.73

d)

All that apply:

  • B. It could be that therapist empathy causes patient satisfaction. This is a reasonable direction of causality, as patients are likely to appreciate the empathy.
  • C. It could be that patient satisfaction causes therapist empathy. This is a reasonable direction of causality, as therapists will likely empathize more with satisfied patients.
  • E. It could be that inherent attitude influences empathy and satisfaction. This is a reasonable direction of causality, as some people are naturally happier than others.